OpenStudy (anonymous):

8 5/6 - 2 1/12 simpliest form

OpenStudy (anonymous):

\[20/3 - 1/10\]

OpenStudy (anonymous):

firstly, write each term as a simple fraction, i.e 8 5/6= ((6*8)+5)/6=53/6 then find a common denominator of the two numbers , in this case , 12, and write each term with this as the denominator. , so 53/6=106/12 now add the two terms, giving a fraction n/12 . now simplify this by dividing top and bottom by common factors as much as possible. then separate if you wish into a mixed fraction. you should get an answer of n/4.